A professional hockey player played 20 seasons. The following table shows the number of games in which he played during each of his 20 seasons. Complete parts (a) through (e) below. a. Obtain and interpret the quartiles. 31 (Type integers or decimals.)
Answer #1

Solution

Arranging number of games in ascending order:

45,48, 64,67,69,69,71,72,73,75,75,76,77,77,77,78,78,78,80,82 ; n = 20

Q = (n+1)/4

Q1 = (21/4) = 5.25 = 69

Q2 = (21/2) = 10.5 = 75

Q3 = 3(21/2) = 15.75 = 77*0.25+78*0.75 = 77.75